OpenStudy (yellowlegoguy99):

Match each percent proportion with the percent problem it represents. Terms: 45/100 = 20/x 20/100 = 45/x x/20 = 45/100 45/20 = x/100 Definitions: 45 is what percent of 20? 20% of what number is 45? 20 is 45% of what number? What is 45% of 20?

OpenStudy (yellowlegoguy99):

I don't understand this either...@skullpatrol

OpenStudy (yellowlegoguy99):

@skullpatrol

OpenStudy (austinl):

These are a little bit difficult to answer in the sense of guiding. I understand these, but I am having problems thinking of ways to put the explanation in coherent terms to make it so you can actually understand me.

OpenStudy (skullpatrol):

Let @austinL try @yellowlegoguy99

OpenStudy (yellowlegoguy99):

...?

OpenStudy (austinl):

Ok, out of the top group, lets start with the bottom term. \(\displaystyle \frac{45}{20}=\frac{x}{100}\) Any number over over 100 means that it can automatically be converted in to a percentage. So for this one, 45 is exactly opposite of the x. So therefore we would be solving for, 45 is x percent of 20. Does this make sense?

OpenStudy (yellowlegoguy99):

Yeah

OpenStudy (yellowlegoguy99):

So 45/20 = x/100 = 45 is what percent of 20?

OpenStudy (yellowlegoguy99):

@austinL

OpenStudy (austinl):

Correct.

OpenStudy (austinl):

Now look at the top term, what do you think it is?